# ARDuP: Active Region Video Diffusion for Universal Policies

### v1

@@ I Introduction
One promising direction is jointly using video and text descriptions to define a generalized state and action space. In these scenarios, a video generator is employed as a planner, which produces a sequential trajectory of frames as states given a text description of the immediate goal and an initial visual representation of the environment. Once the trajectory is generated, a policy conditioned on this trajectory (sequence of frames) is learned to infer the action taken between adjacent frames. The intuition is that using videos to represent the state space enables greater generalization across various tasks and environments. Recently, there has been considerable progress made in this field, with notable works [1, 2, 3] demonstrating success in tasks such as robot navigation and manipulation. However, these methods often struggle to solve the task because they generate videos treating all pixels uniformly, often focusing on the wrong areas and neglecting to model pixels that are important for the policy. This can result in errors in generated frames, such as those seen on the top right of Figure 1. The robot picks up the wrong block (purple block) and can still reduce the loss in later frames by simply changing the color of the block to match the text description (e.g., changing purple to white).
@@ II-A Video Diffusion Models for Decision Making
Recent advancements in diffusion models have led to their application across various decision-making problems [8, 9, 10, 11]. For instance, Diffuser [12] developed an unconditional diffusion model to generate trajectories comprising states and actions based on joints. As a pioneering approach, UniPi [2] casts the decision-making problem as a text-conditioned video generation problem and uses a video diffusion model as a universal policy. Extending UniPi, [13] leverages dense correspondence for action inference. And [1] incorporates VLMs with video diffusion models to facilitate the generation of long-horizon video plans. Existing works usually use the video diffusion model to generate every pixel indifferently of the videos. In comparison, we propose to emphasize the generation of task-relevant objects, which leads to policy more aligned with the task requirements.
@@ II-B Active Region for Vision and Robotics
Active objects are objects being manipulated by the user and are important to the action being performed [14]. The idea that some regions are more important than others has been investigated in First Person Vision [15] and egocentric activity recognition [16, 17, 18]. In contrast, active regions for video-based policy is under-explored. We emphasize active regions in video generation and align the policy’s attention to active regions based on the task text. Since the ground truth active region is not available in real world, we generate pseudo-active regions from videos via dense point tracking [4] and SAM [5] to eliminate manual labeling.
@@ III-B Latent Unified Predictive Decision Process conditioned on Active Region (LUPDP-AR)
IV Active Region Video Diffusion for Unisersal Policies (ARDuP)
@@ IV-A2 Latent Active Region Diffusion
Diffusion models are flexible, which enables the model to select many possible active regions. This flexibility is important because there are often many ways to complete a task and we want a model that encompasses this uncertainty. For example, in a task described as “put blue blocks into the box”, any unplaced blue block can be identified as a valid active region. This flexibility is particularly beneficial in dynamic environments where multiple paths to task completion may exist.
@@ V Experiments
In the following sections, we evaluate our method on two public robot manipulation datasets, CLIPort [6] and BridgeData v2 [7]. First, we describe the implementation details in Sec. V-A. In Sec. V-B we evaluate the generalization ability of our method across multiple tasks. We provide these results on CLIPort. Sec. V-C demonstrates the transfer ability of our method to real-world complex scenarios. We report results on BridgeData v2. Finally, an ablation study and a comprehensive analysis are provided in Sec. V-D.
@@ V-B Multi-Environment Transfer on CLIPort
Results. In Table I, we show the performance of our method in comparison to established baselines on three unseen tasks in CLIPort. Notably, our method significantly improves the success rate over the UniPi baseline by $21.3\%$, $17.2\%$, and $15.7\%$ for the Place Bowl, Pack Object, and Pack Pair tasks, respectively. This performance is driven by better generation as a result of our active region-conditioned diffusion models and improved execution produced by the latent inverse dynamics module. We visualize generated video plans in Figure 3, highlighting the visual improvements of ARDuPover a counterpart model without active region conditioning. The counterpart is the same latent video diffusion model as ours except for conditioning on the active region. Using an active region results in higher visual quality, especially around objects of interest. For example, our method successfully generates a video that packs the white block, while the counterpart generates frames that either pick the wrong object or fail to correctly generate the arm position.
@@ V-C Real-world Transfer on BridgeData v2
Setup. BridgeData v2 [7] is a large and diverse real-world dataset for robotic manipulation. Bridge’s examples demonstrate a wide variety of skills using a large number of unique objects in diverse environments. This dataset includes 60,096 trajectories with natural language instructions, gathered from 24 different environments using a widely accessible, cost-effective robot. We use 95% of the data for training and the rest for evaluation.

### v2

@@ I Introduction
One promising direction is jointly using video and text descriptions to define a generalized state and action space. In these scenarios, a video generator is employed as a planner, which produces a sequential trajectory of frames as states given a text description of the immediate goal and an initial visual representation of the environment. Once the trajectory is generated, a policy conditioned on this trajectory (sequence of frames) is learned to infer the action taken between adjacent frames. The intuition is that using videos to represent the state space enables greater generalization across various tasks and environments. Recently, there has been considerable progress made in this field, with notable works [1, 2, 3] demonstrating success in tasks such as robot navigation and manipulation [4]. However, these methods often struggle to solve the task because they generate videos treating all pixels uniformly, often focusing on the wrong areas and neglecting to model pixels that are important for the policy. This can result in errors in generated frames, such as those seen on the top right of Figure 1. The robot picks up the wrong block (purple block) and can still reduce the loss in later frames by simply changing the color of the block to match the text description (e.g., changing purple to white).
@@ II-A Video Diffusion Models for Decision Making
Recent advancements in diffusion models have led to their application across various decision-making problems [9, 10]. For instance, Diffuser [11] developed an unconditional diffusion model to generate trajectories comprising states and actions based on joints. As a pioneering approach, UniPi [2] casts the decision-making problem as a text-conditioned video generation problem and uses a video diffusion model as a universal policy. Extending UniPi, [12] leverages dense correspondence for action inference. And [1] incorporates VLMs with video diffusion models to facilitate the generation of long-horizon video plans. Existing works usually use the video diffusion model to generate every pixel indifferently of the videos. In comparison, we propose to emphasize the generation of task-relevant objects, which leads to policy more aligned with the task requirements.
@@ II-B Active Region for Vision and Robotics
Active objects are objects being manipulated by the user and are important to the action being performed [13]. The idea that some regions are more important than others has been investigated in First Person Vision [14] and egocentric activity recognition [15, 16]. In contrast, active regions for video-based policy are under-explored. We emphasize active regions in video generation and align the policy’s attention to active regions based on the task text. Since the ground truth active region is not available in real world, we generate pseudo-active regions from videos via dense point tracking and SAM to eliminate manual labeling [5, 17, 18, 19, 20, 21, 6].
@@ III-B Latent Unified Predictive Decision Process conditioned on Active Region (LUPDP-AR)
IV Active Region Video Diffusion for Universal Policies (ARDuP)
@@ IV-A2 Latent Active Region Diffusion
Diffusion models offer flexibility, allowing the selection of various active regions, which is crucial for tasks with multiple valid solutions. For example, in a task described as “put blue blocks into the box”, any unplaced blue block can be identified as a valid active region. This flexibility is particularly beneficial in dynamic environments where multiple paths to task completion may exist.
@@ V Experiments
In the following sections, we evaluate our method on two public robot manipulation datasets, CLIPort [7] and BridgeData v2 [8]. First, we describe the implementation details in Sec. V-A. In Sec. V-B we evaluate the generalization ability of our method across multiple tasks. We provide these results on CLIPort. Sec. V-C demonstrates the transfer ability of our method to real-world complex scenarios. We report results on BridgeData v2. Finally, ablation studies are provided in Sec. V-D, and limitations are discussed in Sec. V-E.
@@ V-B Multi-Environment Transfer on CLIPort
Results. In Table I, we show the performance of our method compared with established baselines on three unseen tasks in CLIPort. Notably, our method significantly improves the success rate over the UniPi baseline by $21.3\%$, $17.2\%$, and $15.7\%$ for the Place Bowl, Pack Object, and Pack Pair tasks, resp. This performance is driven by better generation due to our active region-conditioned diffusion models and improved execution produced by the latent inverse dynamics module. We visualize generated video plans in Figure 3, highlighting the visual improvements of ARDuP over a counterpart model without active region conditioning. The counterpart is the same latent video diffusion model as ours except for conditioning on the active region. Using an active region results in higher visual quality, especially around objects of interest. For example, our method successfully generates a video that packs the white block, while the counterpart generates frames that either pick the wrong object or fail to generate the arm position correctly.
@@ V-C Real-world Transfer on BridgeData v2
Setup. BridgeData v2 [8] is a large and diverse real-world dataset for robotic manipulation. Bridge’s examples show a wide variety of skills using various unique objects in diverse environments. This dataset includes 60,096 trajectories with natural language instructions, gathered from 24 environments with a widely accessible, cost-effective robot. We use 95% of the data for training and the rest for evaluation.
@@ V-D2 Effect of Active Region Quality on Performance
V-E Limitations and Future work
One limitation of our approach is that pseudo-active region quality depends on Co-Tracker, which may struggle with long-horizon tasks or complex backgrounds. Advanced dense point tracking models could improve pseudo label quality, and noisy label learning techniques might stabilize training. Alternatively, text-guided segmentation could be used, where the vision-language model identifies the target object based on the task, followed by segmentation using tools like Segment Anything. We leave these as future work.
@@ VI Conclusion
Acknowledgements This work was partially supported by NSF CAREER Award (#2238769) to AS.